OpenStudy (anonymous):

Can y'all tell me all you know about vectors? I am particularly fuzzy on defining planes using vectors and, in general, figuring out parametric equations. Thanks!

OpenStudy (amistre64):

a plane is defined as a infinite number of vectors that are perpendicular to the normal vector

OpenStudy (amistre64):

a parametric equation is just letting the parts be independant of each other but dependant on a common variable, time usually

OpenStudy (amistre64):

x = 3+4t y = -2-t

OpenStudy (amistre64):

you might wanna have some specific examples to work with tho